OpenStudy (anonymous):

Find all the second partial derivatives. f(x, y) = sin2(mx + ny)

OpenStudy (anonymous):

We have: \[f(x,y)=sin[2(mx+ny)]\] For a function \(f(x,y)\), there exist four possible partial derivatives: \[\frac{\partial }{\partial x}\left[\frac{\partial f}{\partial x}\right]=f_{xx}=a\] \[\frac{\partial }{\partial x}\left[\frac{\partial f}{\partial y}\right]=f_{xy}=b\] \[\frac{\partial }{\partial y}\left[\frac{\partial f}{\partial x}\right]=f_{yx}=c\] \[\frac{\partial }{\partial y}\left[\frac{\partial f}{\partial y}\right]=f_{yy}=d\] We will calculate each!

OpenStudy (anonymous):

Good so far? Makes sense?

OpenStudy (anonymous):

Lets start: \[a=\frac{\partial}{\partial x}\left[\frac{\partial f}{\partial x}\right]=\frac{\partial}{\partial x}\left[\frac{\partial\bigg(\sin[2(mx+ny)]\bigg)}{\partial x}\right]=\frac{\partial}{\partial x}\left[\frac{\partial\Big(\sin(2mx+2ny)\Big)}{\partial x}\right]\] \[=\frac{\partial}{\partial x}\left[2m\cos(2mx+2ny)\right]=-4m^2\sin(2mx+2ny)\]

OpenStudy (anonymous):

Cool?